OpenStudy (anonymous):

ok -5x + 3y = 6;(-3,-8)

OpenStudy (anonymous):

I'm guessing this is a parallel line question. So first rearrange the equation. So you get 3y=6+5x. Now divide all by 3 so you get y=2+(5/3)x. So the new parallel line must ALSO have the slope (5/3). So the new line must have the form y=(5/3)x+b. Now all you gotta do from here is see if you're given a point that you know will be contained in this new line. Which is (-3,-8). So plug in y=-8 and x=-3, and find b.

OpenStudy (anonymous):

thank you soo much!